OpenStudy (anonymous):

Daniel is looking over some data regarding the temperatures of countries and how the temperature relates to the number of mosquitoes. The data are represented by the exponential function f(x) = 2x, where x is the number of mosquitoes. Explain how he can convert this equation to a logarithmic function when the temperature is 64 degrees. Can someone help me and explain this to me?

OpenStudy (welshfella):

f(x) = 2^x ln f(x) = ln 2^x ln f(x) = x ln 2

OpenStudy (welshfella):

ln 64 = x ln 2

OpenStudy (anonymous):

so I plugged 64 in for F(x)? do I divide both sides after that?

OpenStudy (welshfella):

yes divide both sides by ln2 to get x

OpenStudy (anonymous):

okay, thank you :)

OpenStudy (welshfella):

yw